# 3DEnhancer: Consistent Multi-View Diffusion for 3D Enhancement

### v1

@@ 4.2 Comparisons
Figure 5: Qualitative comparisons of enhancing 3D reconstruction given generated multi-view images on the in-the-wild dataset. Due to the low quality and view-inconsistency of outputs from multi-view models, the reconstructed 3D models often suffer from issues such as blurry textures, like the skull’s teeth in the first row, or artifacts such as ghosting hands in the second Mario example. Existing enhancement methods fail to correct texture artifacts. In contrast, our method produces both geometrically accurate and visually appealing results. (Zoom-in for best view)
Figure 6: Low-resolution 3D Gaussians optimization with 3DEnhancer. (Zoom-in for best view)
Enhancing 3D Reconstruction. In this section, we present 3D reconstruction comparisons based on rendering views from 3D Gaussians generated by LGM [58]. Quantitative comparisons are shown in Tab. 3. The proposed method outperforms previous approaches in terms of 3D reconstruction. For qualitative evaluation, we visualize the results of two video enhancement models, RealBasicVSR and Upscale-A-Video. As shown in Fig. 5, these baselines suffer from a lack of multi-view consistency, leading to misalignment, such as the misalignment of the teeth in the first skull example and the ghosting in the example of Mario’s hand. In contrast, our model maintains consistency and produces high-quality texture details. We further demonstrate our approach can optimize a coarse differentiable representation. As shown in Fig. 6, our method is capable of refining low-resolution Gaussians [52]. More details and results of refining coarse Gaussians are provided in the Appendix Sec. D.2.
@@ 4.3 Ablation Study
Figure 7: Effectiveness of cross-view modules.
Figure 8: Comparisons of enhancing multiview images with and without epipolar aggregation. The red line denotes the epipolar line corresponding to the circled area, while the dotted arrow indicates the target location of tokens propagated from one view to another. (Zoom-in for best view)
Effectiveness of Noise Level. As shown in Fig 1, our model can generate diverse textures by adjusting noise levels. Low noise levels generally result in outputs with blurred details, while high noise levels produce sharper, more detailed textures. However, high noise levels may also reduce the fidelity of the input images.
@@ 5 Conclusion
In this appendix, we provide additional discussions and results to supplement the main paper. In Sec. A, we present more architecture and design details of our 3DEnhancer. In Sec. B, we provide detailed information about our training dataset, including the augmentation pipeline and illustrative examples. Sec. C highlights some interesting findings related to inference. Comprehensive results are presented in Sec. D to further demonstrate our performance. Notably, we also include a demo video (Sec. D.4) to showcase rendering results for 3D reconstruction enhancement.
@@ C.1 Multi-View Editing
Benefiting from our comprehensive augmentation pipeline and the robust view-consistent DiT Block, we observe an interesting fact: our method is capable of generating detailed and consistent textures even from extremely coarse or corrupted multi-view inputs. This enables our approach to modify multi-view images in two distinct ways: 1. Applying a black mask to the region designated for editing and modifying the text prompt to generate the target multi-view images. 2. Adjusting the inference noise level, where higher noise levels produce more diverse outputs. Using the edited multi-view images, we can subsequently modify the reconstructed 3D representations. An example of editing 3D Gaussians generated by LGM [58] through modifying its multi-view input is shown in Fig. 10.
@@ D.2 Results of Optimizing 3D Gaussians
Table 7: Quantitative comparison of optimizing low-resolution Gaussians. The best results are highlighted in bold.
@@
D.3 More Comparisons
In this section, we introduce another baseline from the multi-view image upscale module in Unique3D [72]. This baseline finetunes ControlNet-Tile [83] to enhance RGB views. While the module can sharpen some textures, it struggles to recover inconsistent or corrupted areas in multi-view images. Our method outperforms Unique3D’s MV Upscale both quantitatively and qualitatively. The quantitative comparison between Unique3D’s MV Upscale and our method is presented in Tab. 8. Additionally, we provide more visual comparisons of our method with all other baselines, including RealESRGAN [69], StableSR [64], Unique3D’s MV Upscale [72], RealBasicVSR [7], and Upscale-a-Video [91]. Fig. 14 and Fig. 15 showcase the visual comparisons of multi-view enhancement on synthetic and in-the-wild datasets, respectively.
Table 8: Quantitative comparison of enhancing multi-view synthesis on the Objaverse synthetic dataset with Unique3D’s MV Upscale module. Our method demonstrates clear advantages in restoration fidelity, as measured by PSNR, SSIM, and LPIPS. While applying color correction improves the output of Unique3D’s MV Upscale module, it has minimal impact on our results when noise level is set to 0, highlighting our method’s robustness against global color scheme shift issues.

### v2

@@ 4.2 Comparisons
Figure 5: Qualitative comparisons of enhancing 3D reconstruction given generated multi-view images on the in-the-wild dataset. Multi-view models produce low-quality, view-inconsistent outputs, leading to flawed 3D reconstructions. Existing methods fail to correct texture artifacts, while our method produces both geometrically accurate and visually appealing results.
Figure 6: Low-resolution GS optimization with 3DEnhancer.
Enhancing 3D Reconstruction. In this section, we present 3D reconstruction comparisons based on rendering views from 3DGS generated by LGM [60]. Quantitative comparisons are shown in Tab. 3. Our method outperforms previous approaches in terms of 3D reconstruction. For qualitative evaluation, we visualize the results of two video enhancement models, RealBasicVSR and Upscale-A-Video. As shown in Fig. 5, these baselines suffer from a lack of multi-view consistency, leading to misalignment, such as the misalignment of the teeth in the first skull example and the ghosting in the example of Mario’s hand. In contrast, our model maintains consistency and produces high-quality texture details. We further demonstrate our approach can optimize coarse differentiable representations. As shown in Fig. 6, our method is capable of refining low-resolution Gaussians [54]. More details and results of refining coarse Gaussians are provided in the Appendix Sec. D.2.
@@ 4.3 Ablation Study
Figure 8: Comparisons of enhancing multi-view images with and without epipolar aggregation. The red line denotes the epipolar line corresponding to the circled area, while the dotted arrow indicates the corresponding area from one view to another.
Effectiveness of Noise Level. As shown in Fig. 1, our model can generate diverse textures by adjusting noise levels. Low noise levels generally result in outputs with blurred details, while high noise levels produce sharper, more detailed textures. However, high noise levels may also reduce the fidelity of the input images.
@@ 5 Conclusion
Acknowledgement. This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s). It is also supported by Singapore MOE AcRF Tier 2 (MOE-T2EP20221-0011) and the National Research Foundation, Singapore, under its NRF Fellowship Award (NRF-NRFF16-2024-0003).
@@
In this appendix, we provide additional discussions and results to supplement the main paper. In Sec. A, we present more architecture and design details of our 3DEnhancer. In Sec. B, we provide detailed information about our training dataset, including the augmentation pipeline and illustrative examples. Sec. C highlights some interesting findings related to inference. More results and comparisons are presented in Sec. D to further demonstrate our performance. We also include a demo video (Sec. D.6) to showcase rendering results for 3D reconstruction enhancement.
@@ C.1 Multi-View Editing
Benefiting from our comprehensive augmentation pipeline and the robust view-consistent DiT Block, we observe an interesting fact: our method is capable of generating detailed and consistent textures even from extremely coarse or corrupted multi-view inputs. As shown in Fig. 10, our method effectively handles various challenging cases, including multi-views with (a) extremely blurred textures, (b) masked or missing parts, and (c) significant noise.
Figure 10: Examples of handling extremely coarse inputs with 3DEnhancer.
This enables our approach to modify multi-view images in two distinct ways: 1. Applying a black mask to the region designated for editing and modifying the text prompt to generate the target multi-view images. 2. Adjusting the inference noise level, where higher noise levels produce more diverse outputs. Using the edited multi-view images, we can subsequently modify the reconstructed 3D representations. An example of editing 3D Gaussians generated by LGM [60] through modifying its multi-view input is shown in Fig. 11.
@@ D.2 Results of Optimizing 3D Gaussians
Table 7: Quantitative comparisons of optimizing low-resolution Gaussians. The best results are highlighted in bold.
@@
D.3 Results of Generalization to Real-World Objects
We test our model on the constructed OmniObject3D dataset [76], which provides realistic 3D object scans, and also on complex, richly textured objects from Polycam. Backgrounds are removed as needed using BiRefNet [90]. As shown in Fig. 15, our model effectively enhances real-world objects.
Figure 15: Examples of handling complex real-world objects with 3DEnhancer. Our method generates rich textures on realistic objects.
D.4 Results of Further Fine-tuning Upscale-A-Video
Our work aims to provide a generic framework for 3D object enhancement, supporting enhancing (I) sparse multi-view images from large angles for multi-view reconstruction networks (e.g., LGM [60]), and (II) coarse 3D model via per-instance optimization. Existing video diffusion models, e.g., Upscale-A-Video [95], mainly rely on temporal attention for consistency. They are designed for handling adjacent video frames with minimal spatial variations, without considering camera pose. Thus, they struggle to establish multi-view correspondences in case (I), where input views vary significantly, leading to suboptimal results. Additionally, due to the huge GPU memory cost of video diffusion models, they also cannot handle dense 360° views simultaneously, typically working with short video sequences (e.g., 8 frames for Upscale-A-Video), limiting their performance in case (II) as well. Thus, this study is crucial to explore new and effective modules of sparse multi-view attention for 3D enhancement, using a pose-aware encoder and an epipolar aggregation mechanism, which together achieve superior results in both (I) and (II) (see Tabs. 1-3 and Figs. 3 and 4). All baseline methods in the main paper are fine-tuned on the Objaverse dataset. We further fine-tune Upscale-A-Video with our proposed data augmentation. The results in the Tab. 8 and Fig. 16 show that our method still outperforms the video-based Upscale-A-Video, further supporting our discussion here.
Table 8: Quantitative comparisons with fine-tuned Upscale-A-Video (UAV) on synthetic Objaverse multi-view images and Low-Resolution (LR) Gaussians.
Method Synthetic Objaverse Low-Resolution Gaussians
[…]
UAV (main paper) 25.57 0.8937 0.1153 26.20 0.9184 0.0928
UAV (further fine-tuned) 26.14(+0.57) 0.9086(+0.0149) 0.0996(-0.0157) 26.69 (+0.49) 0.9197 (+0.0013) 0.0850 (-0.0078)
Ours 27.53 0.9265 0.0626 27.54 0.9337 0.0756
Figure 16: Qualitative comparisons with fine-tuned Upscale-A-Video (UAV) on synthetic Objaverse multi-view images and Low-Resolution (LR) Gaussians. With additional fine-tuning using our augmentations, Upscale-A-Video reduces inconsistent artifacts outside the object area. Our method still shows superior generative capabilities.
D.5 More Comparisons
In this section, we introduce another baseline from the multi-view image upscale module in Unique3D [74]. This baseline fine-tunes ControlNet-Tile [86] to enhance RGB views. While the module can sharpen some textures, it struggles to recover inconsistent or corrupted areas in multi-view images. Our method outperforms Unique3D’s MV Upscale both quantitatively and qualitatively. The quantitative comparisons between Unique3D’s MV Upscale and our method is presented in Tab. 9. Additionally, we provide more visual comparisons of our method with all other baselines, including RealESRGAN [71], StableSR [66], Unique3D’s MV Upscale [74], RealBasicVSR [7], and Upscale-a-Video [95]. Fig. 17 and Fig. 18 showcase the visual comparisons of multi-view enhancement on synthetic and in-the-wild datasets, respectively.
Table 9: Quantitative comparisons of enhancing multi-view synthesis on the Objaverse synthetic dataset with Unique3D’s MV Upscale module. Our method demonstrates clear advantages in restoration fidelity, as measured by PSNR, SSIM, and LPIPS. While applying color correction improves the output of Unique3D’s MV Upscale module, it has minimal impact on our results when noise level is set to 0, highlighting our method’s robustness against global color scheme shift issues.